OpenStudy (vegan555):

A man wandering in the desert walks 2.6 miles in the direction S 34° W. He then turns 90° and walks 3.3 miles in the direction N 56° W. At that time, how far is he from his starting point, and what is his bearing from his starting point?

Mehek (mehek14):

The picture would be a right triangle - his two walking paths are the lags and the the hypotenuse is his total distance from the starting point. Use the Pythagorean Theorem.

OpenStudy (vegan555):

so that would be the miles from his starting point? I got 4.2

OpenStudy (trojanpoem):

N W is north west ? S W is south west ?

OpenStudy (vegan555):

yes

OpenStudy (trojanpoem):



OpenStudy (trojanpoem):

He went from C to A and from A to B so the starting point mentioned is A to B so AB= rootof((2.6)^2 + (3.3)^2) = some number.

OpenStudy (vegan555):

4.2

OpenStudy (trojanpoem):

Yeah.

OpenStudy (vegan555):

so that's the miles

OpenStudy (vegan555):

so how do I find the bearings?

OpenStudy (trojanpoem):

Bearing mean ?

OpenStudy (vegan555):

his bearing from his starting point

OpenStudy (trojanpoem):

Direction ?

OpenStudy (vegan555):

the number of degrees from his starting point I guess

OpenStudy (vegan555):

this is really confusing to me sorry

OpenStudy (trojanpoem):

You can use the sin or cos or tan

OpenStudy (vegan555):

would it be cos56 degrees?

OpenStudy (trojanpoem):

tan(a) = 3.3/2.6 so a =51 46

OpenStudy (vegan555):

wouldn't it be tan 2.3/4.2?

OpenStudy (vegan555):

oh wait nevermind sorry

OpenStudy (trojanpoem):

tan = BC/AC Cos = AC/AB Sin = BC/AB

OpenStudy (vegan555):

ok so 2.3/3.3= .696969, so then do I do tan^-1(.696969)? I got 34.875

OpenStudy (trojanpoem):

2.6/3.3

OpenStudy (vegan555):

oh shoot ok sorry thank you lol

OpenStudy (vegan555):

ok so I got 38.2

OpenStudy (trojanpoem):

Sorry 3.3/2.6

OpenStudy (trojanpoem):

I didn't look at the triangle :ops

OpenStudy (vegan555):

ooh ok

OpenStudy (vegan555):

well tan is opp/adj so wouldn't it be 2.6/3.3?

OpenStudy (trojanpoem):

I think what confuses you is how to get tan sin and cos right ?

OpenStudy (vegan555):

yes lol

OpenStudy (trojanpoem):



OpenStudy (trojanpoem):

The sin is the side in front of it (x) / the side which is in front of the 90 degree

OpenStudy (trojanpoem):

the cos is the side (V)/ the side in front of the 90

OpenStudy (trojanpoem):

tan is the side (x) (front) / the side (v)

OpenStudy (vegan555):

ok I'm sorry, I thought tan was v/x not x/v. SOH CAH TOA lol. Tan is opp/adj.

OpenStudy (trojanpoem):

Saying opp/adj was confusing m in the past LAWL

OpenStudy (trojanpoem):

N.m , good luck.